OpenStudy (anonymous):

lim x→0 (cos 5x tan 5x)/x

OpenStudy (anonymous):

\[\tan5x=\frac{\sin 5x}{\cos5x}\] so \[\cos5x\tan5x=\sin5x\] \[\lim_{x\to0}\frac{\sin5x}{x}=\lim_{x\to0}\frac{\sin5x}{5x}\cdot5=\cdots \]

OpenStudy (anonymous):

Ok I follow the logic thanks for the help

OpenStudy (anonymous):

yw